Operations Report – April 20th

Crew 192 Operations Report 20Apr2018

SOL 12

Name of person filing report: Richard Blakeman

Non-nominal systems: Upper deck faucet, hot water now has to be turned off when not in use because it leaks too much, causing water loss and constant noise from the pump.

Generator (hours run): 12 hours

Solar ‒ SOC 87%

Diesel – 75% fuel

Propane – 60% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 25% gas left in one can.

Water (trailer) – Empty water tank secure on Trailer.
Water (static) – 550 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133825.0

Toilet tank emptied: Yes. Last emptied today, April 20.

ATV’s Used: Blue ATVs 1, 2, 3, and 4

Oil Added? No but checked and OK last time road.

ATV Fuel Used: 1, 2, and 3 at 75%, 4 at 80%

# Hours the ATVs were Used today: 6

Notes on ATVs: ATVs fine, oil checked.

Deimos rover used: No
Hours: 122.3
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Opportunity rover used: No
Hours: 30.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 27.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Cleaned the hab from top to bottom to prepare for incoming crew.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, two one-piece suits used, radios working fine. Comms headpieces all cleaned and prepared for incoming crew.

Summary of Hab operations: Cleaning and organizing for Crew 193.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: Last cleaning check.

Summary of RAM operations: None

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: None

Operations Report – April 19th

Crew 192 Operations Report 19Apr2018

SOL 11

Name of person filing report: Richard Blakeman

Non-nominal systems: Hot water faucet in upper deck of hab, cleaned out water pump for water tank.

Generator (hours run): 12 hours

Solar ‒ SOC 87%

Diesel – 75% fuel

Propane – 61% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – All ATVs filled 100%, 25% left in one can.

Water (trailer) – Empty water tank secure ob Trailer.
Water (static) – 690 gallons (two external tanks)
Trailer to Static Pump used – Yes
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133755.8

Toilet tank emptied: No. Last emptied April 16.

ATV’s Used: None

Oil Added? No but checked and OK last time road.

ATV Fuel Used: 0% today

# Hours the ATVs were Used today: 0

Notes on ATVs: ATVs fine, oil checked, gas full.
Assumption on percentage …untenable to check
Deimos rover used: No
Hours: 121.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: N/A
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: N/A
Currently charging: Yes

Curiosity rover used: No
Hours: 27.4
Beginning charge: 100%
Ending charge: N/A
Currently charging: No

HabCar used and why, where? No

General notes and comments: A great EVA was followed by cleaning and preparing the hab for Crew 193.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, three one-piece suits used, radios working fine. Inventoried and cleaned 30 ear pieces not including 5 our crew has.

Summary of Hab operations: Cleaning and organizing for Crew 193.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: Thoroughly cleaned and prepared for next crew.

Summary of RAM operations: Trailer water tank moved out of RAM and back onto trailer and secured. Doctor requests water test kit before use. Cleaned, swept and mopped RAM…secured doors.

Summary of health and safety issues: Crew was very tired after working on water tank, so afternoon EVA was changed to an engineering EVA.

Questions, concerns, and requests to Mission Support: Fuel for ATVs

Operations Report – April 18th

Crew 192 Operations Report 18Apr2018

SOL 10

Name of person filing report: Victoria Varone

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 66%

Diesel – 75% fuel

Propane – 62% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – All ATVs filled 100%, 25% left in one can, requesting gas refill.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 690 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133669.6

Toilet tank emptied: No. Last emptied April 16.

ATV’s Used: None

Oil Added? No but checked and OK.

ATV Fuel Used: 0% today

# Hours the ATVs were Used today: 0

Notes on ATVs: ATVs fine, oil checked, gas filled yesterday.

Deimos rover used: Yes (out of commission per Shannon)
Hours: 121.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: 68%
Currently charging: Yes

Opportunity rover used: yes
Hours: 29.8
Beginning charge: 100%
Ending charge: 81%
Currently charging: Yes

Curiosity rover used: Yes
Hours: 27.4
Beginning charge: 100%
Ending charge: 67%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Some great EVAs on a beautiful day.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, three one-piece suits used, radios working fine.

Summary of Hab operations: Cooked, cleaned, and organized the hab.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water tank remains in the RAM to air out after being cleaned for the final time yesterday, will be checked tomorrow.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: – Hab upper deck faucet hot water shut due to increasing leaks.

– Fridge got hot last night despite being closed and plugged in properly, was unplugged and plugged back and is working fine again.

– Steps in front of rear airlock were secured again after become a bit loose after the first fix.

Operations Report – April 17th

Crew 192 Operations Report 17Apr2018

SOL 9

Name of person filing report: Victoria Varone

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 87%

Diesel – 75% fuel

Propane – 62% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – All ATVs filled 100%, 25% left in one can, requesting gas refill.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 730 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133616.7

Toilet tank emptied: No. Emptied yesterday, April 16.

ATV’s Used: Red ATV, Blue ATV’s #1, #2, and #3

Oil Added? No but checked and OK.

ATV Fuel Used: 20%

# Hours the ATVs were Used today: 6

Notes on ATVs: All ATVs working well, oil checked, gas filled today.

Deimos rover used: No
Hours: 120
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 31.2 (corrected)
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.9
Beginning charge: 91%
Ending charge: 91%
Currently charging: Yes

HabCar used and why, where? No

General notes and comments: Crew is doing well, trying to make the best use of time for EVAs in the midst of constant strong winds.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, three two-piece suits used, radios working fine. Suit #4 charging port is faulty.

Summary of Hab operations: Cleaned and organized hab.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water was cleaned for final time today, remains in the RAM while strong winds persist.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: Doctor Narayanamoorthi recommends biological and chemical testing of water in trailer water tank before clearance for crew use.

Faucet in upper deck of hab was fixed as best we could but it continues to slowly leak.

Operations Report – April 16th

Crew 192 Operations Report 16Apr2018

SOL 8

Name of person filing report: Richard Blakeman

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 78%

Diesel – 70% fuel

Propane – 63% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Approximately 4 gallons left in one and the other is empty

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 780 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133570.4

Toilet tank emptied: Yes. Emptied today, April 16.

ATV’s Used: None

Oil Added? No but checked and OK.

ATV Fuel Used: 5%

# Hours the ATVs were Used today: 0.5

Notes on ATVs: All working well and ready for use, gas filled up yesterday.

Deimos rover used: Yes
Hours: 120
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 38.2
Beginning charge: 100%
Ending charge: 98%
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.9
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well, though the bad weather is hampering our EVAs and astronomical activities.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, 3 two-piece suits used, radios working fine.

Summary of Hab operations: Finished removing hab skirt as requested, kept a close eye on all hab systems as strong winds blew most of the day.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water remains in the RAM for continued cleaning of embedded sediment.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: Requesting trash pickup.

Operations Report – April 15th

Crew 192 Operations Report 15Apr2018

SOL 7

Name of person filing report: Richard Blakeman

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 84%

Diesel – 75% fuel

Propane – 64% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Approximately 4 gallons left in one and the other is empty

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 815 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133514.4

Toilet tank emptied: Yes. Emptied today, 15 April.

ATV’s Used: Blue ATVs #1, #2

Oil Added? No but checked and OK.

ATV Fuel Used: 5%

# Hours the ATVs were Used today: 0.5

Notes on ATVs: All working well and ready for use, gas filled up today, all at 100%.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 38.2
Beginning charge: 100%
Ending charge: 98%
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.9
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: A quiet day at the hab.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, two one-piece suits and two two-piece suits used. Radios working fine. Spacesuit 9 ventilation wasn’t working yesterday, but worked today. May have just needed a charge/charging port may not have been inserted all the way. All good.

Summary of Hab operations: Stand-down day, crew relaxed around the hab. Also took the opportunity to perform some training. See commander’s report.

Summary of GreenHab operations: Plants watered twice.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water tank prepped for additional cleaning in the RAM.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: None

Operations Report – April 14th

Crew 192 Operations Report 12Apr2018

SOL 6

Name of person filing report: Victoria Varone

Non-nominal systems: Internet. WiFi was down for most of the day, every member of the crew was getting a “WiFi Connected, but No Internet” message. It’s holding now after being reset but we will keep monitoring it.

Generator (hours run): 12 hours

Solar ‒ SOC 84%

Diesel – 75% fuel

Propane – 66% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Shannon told us we were not to worry about this gas. We have no knowledge of the gas or that 5 gallon container.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 850 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133477.4

Toilet tank emptied: No. Last emptied April 11, 2018.

ATV’s Used: Red ATV, Blue ATVs #1, #2, and #2

Oil Added? No but checked and OK.

ATV Fuel Used: 20%

# Hours the ATVs were Used today: 6

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: Yes
Hours: 25.9
Beginning charge: 100%
Ending charge: 91%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well and having a great time going on EVAs and cooking together.

Summary of internet: 697MB at 2040.

Summary of suits and radios: Suits all charging, 3 one-piece suits and 2 two-piece suits used. Radios working fine.

Summary of Hab operations: Two successful EVAs, inspected and tested Musk Observatory (first day of acceptable weather), lots of cooking.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None

Summary of RAM operations: The trailer water tank remains in the RAM for additional cleaning.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: Spacesuit 9’s ventilation fan is not working. It worked during our first EVA but is not charging.

Operations Report – April 13th

Crew 192 Operations Report 12Apr2018

SOL 5

Name of person filing report: Richard Blakeman

Non-nominal systems: 1. Refrigerator (resolved); 2. Kitchen sink faucet

Notes on non-nominal systems: The refrigerator appeared to have stopped working overnight. Outlet was inspected, fridge unplugged and plugged back in, has been working since. Will continue to monitor.

Sink is being worked on right now.

Generator (hours run): 13 hours

Solar ‒ SOC 98%

Diesel – 75% fuel

Propane – 66% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Shannon told us we were not to worry about this gas. We have no knowledge of the gas or that 5 gallon container.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 925 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133428.9

Toilet tank emptied: No. Last emptied April 11, 2018.

ATV’s Used: Blue ATVs #1 and #3

Oil Added? No but checked and OK.

ATV Fuel Used: 10%

# Hours the ATVs were Used today: 3

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, 2 one piece suits used. Radios working fine.

Summary of Hab operations: Cleaned and organized hab, including ceiling lights in the upper deck and various bins and shelves in lower deck. Organized food cabinets and cleaned additional kitchen appliances and pans. Fixed kitchen sink faucet.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None

Summary of RAM operations: The trailer water tank remains in the RAM for additional cleaning.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support:

– Regarding use of tables in the lower deck: There is a 3 foot wooden table in the spacesuit room and two 5 foot plastic tables in the lower deck. Requesting permission to swap the wooden table with one of the longer plastic ones so we can accommodate all one-piece EVA suits in the spacesuit room.

– Request permission to change out water filter that was delivered earlier today.

– Reviewed possible fixes for the skirt of the HAB as requested. Suggest using sheet metal screws and the electric drill to fasten skirting in place. Respectfully request permission to perform this repair via engineering EVA.

– There’s a picture attached of the base of a small tv screen/computer monitor that is in the hab, but does not seem to have the proper cords available for use. Does anyone know if the cords are available anywhere in the hab site?

– The laptop labeled “Hab” in the upper deck also does not have the proper cords to check for power. Please advise.

Thank you for removing the unneeded items we left outside. There are a few additional items we put outside for pickup.

Operations Report – 12 April 2018

SOL 4

Name of person filing report: Joe Dituri

Non-nominal systems: 1. The main hab door broke during the wind.
2. The circuit breaker tripped.

Notes on non-nominal systems: During severe winds again today, the door swung open HARD and almost came off its hinge. We’ve performed a sturdy and strong repair that is holding beautifully so far.

After it became very cold in the hab, we checked the circuit breaker and noticed it had tripped. It was reset.

Generator (hours run): 13 hours

Solar ‒ SOC 98%

Diesel – 75% fuel

Propane – 66% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 960 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133383.5

Toilet tank emptied: Yes. Last emptied April 11, 2018.

ATV’s Used: Blue ATVs #1, #2, and #3

Oil Added? No But checked and OK.

ATV Fuel Used: 10%

# Hours the ATVs were Used today: 2

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Spirit rover used: No
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well, we had to stay on alert today for the strong winds. Will remain alert for continued winds tomorrow.

Summary of internet: 198MB at 1930.

Summary of suits and radios: Suits all charging, 4 one-piece suits and 1 two piece suit used today, radios working fine.

Summary of Hab operations: Repaired main hab door latch and wooden rail after strong winds blew the door all the way out, broke the latch, and knocked the rail off the stairs. Cleaned and decontaminated the trailer water tank in the RAM. The tank requires further cleaning.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None

Summary of RAM operations: The trailer water tank was cleaned in the RAM.

Operations Report – April 11th

Crew 192 Operations Report 11Apr2018

SOL 3

Name of person filing report: Joe Dituri

Non-nominal systems: Structural – the main hab/EVA door.

Notes on non-nominal systems: The door swung open during the severe wind storm and knocked the top piece of the railing off. We secured the door as best we could and it swung open again. We secured it better and it’s holding for now.

Generator (hours run): 13 hours

Solar ‒ SOC 98%

Diesel – 75% fuel

Propane – 67% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Shannon told us we were not to worry about this gas,

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 994 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133345.0

Toilet tank emptied: No. Last emptied April 9.

ATV’s Used: Red ATV, Blue ATV #2

Oil Added? No But checked and OK.

ATV Fuel Used: 5% Red, 15% Blue #2

# Hours the ATVs were Used today: 2

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Spirit rover used: Yes
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

HabCar used and why, where? No

General notes and comments: Crew is doing well, we had to stay on alert today for the strong winds. Will remain alert for continued winds tomorrow.

Summary of internet: 160MB at 1930.

Summary of suits and radios: Suits all charging, 2 one-piece suits used today, radios working fine.

Summary of Hab operations: Cleaned both decks, moved frig and stove to clean behind, used afternoon EVA to check up around habitat after strong winds, fixed cement step in front of rear airlock as it was not secure and almost fell apart.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: Suits inspected for safety and the Science Dome was cleaned and mopped.

Summary of RAM operations: The trailer water tank that we rolled into the RAM yesterday was examined further so we could devise the best way to clean it over the next few days.

Summary of health and safety issues: In an effort to sweep the upper deck and clean behind the refrigerator, we found a spilled box of matches on the floor near the wall directly between the fridge’s electrical system and propane stove. We removed the matches and swept the area. See Science Report for additional notes.

Questions, concerns, and requests to Mission Support: See Commander’s Report.